# Graph-based Approaches and Functionalities in Retrieval-Augmented Generation: A Comprehensive Survey

### v1

@@ 6.2. Methods Applying the Sequential Pipeline
Keqing (ref:keqing, ) adopts a sequential pipeline to address complex multi-hop question answering. It decomposes complex queries into simpler sub-questions using predefined templates. These sub-questions are then aligned with logical chains on a knowledge graph (KG), guiding the retrieval of candidate entities through multi-hop reasoning. A candidate reasoning module evaluates and selects the correct entities from the KG. Finally, the system generates responses by aggregating reasoning paths and answers to provide an interpretable and precise answer to multi-hop questions.
QA-GNN (ref:qa-gnn, ) integrates LMs and KGs for question answering with a sequential pipeline. The process begins by encoding the QA context including the question and answer options with an LM and retrieving a KG subgraph relevant to the question. A joint ”working graph” is constructed by connecting the QA context node to entities within the KG subgraph. To enhance reasoning, the KG nodes are scored for relevance based on their alignment with the QA context using LM-based scoring. An attention-based GNN performs iterative message passing on this working graph, updating both KG entities and QA context representations. Finally, predictions are made by aggregating the LM output, updated node features, and the working graph representation, achieving accurate and explainable results.
DALK (ref:dalk, ) makes LLMs and KGs benefit from each other through an elaborate sequential pipeline design. At the beginning of the pipeline, LLMs process the scientific corpus and construct two domain-specific KGs from the related literature via pair-wised relation extraction and generative relation extraction, respectively. This part of the pipeline can be seen as LLMs benefit KGs. Then, DALK utilizes a coarse-to-fine sampling method and a retrieval approach to select knowledge from KGs. The selected knowledge as well as the domain questions are sent to LLMs to generate accurate answers, therefore realizing that KGs benefit LLMs.
@@ 6.3. Methods Applying the Loop Pipeline
RGNN-Ret (ref:gnn-ret, ) tackles multi-hop reasoning questions by iteratively decomposing the reasoning process into manageable steps. It incorporates a self-critique mechanism that prompts LLMs to generate sub-questions for each reasoning step. After answering a sub-question with retrieved passages, the LLM evaluates whether the accumulated evidence is sufficient to generate a final answer or if further reasoning steps are required. This iterative process enables the system to dynamically adapt to the complexity of the query. To enhance the retrieval quality across reasoning steps, a Recurrent Graph Neural Network (RGNN) is used. The RGNN integrates the graphs of passages from previous reasoning steps, establishing connections between retrieved passages to ensure that relationships between sub-questions and evidence are maintained. By combining semantic distances and contextual information across steps, the RGNN reduces the impact of incorrect sub-questions and improves retrieval accuracy for supporting passages. The iterative loop continues until the self-critique determines that enough evidence has been collected to answer the question.
ArcaneQA (ref:arcaneqa, ) is designed to address the combinatorial explosion problem commonly encountered in searching for matched subgraphs within large-scale knowledge bases. Unlike traditional approaches that involve an exhaustive search, ArcaneQA narrows its focus to a small set of admissible tokens derived from the relevant subgraph, rather than considering the entire vocabulary. This selective prediction significantly reduces computational complexity. Utilizing an encode-decode architecture built on top of the BERT model, ArcaneQA iteratively predicts additional sequences from the subgraph, allowing for efficient stepwise progression. By breaking down the task of program generation into a sequence of smaller, more manageable decision-making processes, ArcaneQA effectively transforms a daunting search challenge into a tractable series of localized predictions, streamlining the reasoning process while maintaining accuracy.
GRAPH-COT (ref:graphcot, ) provides a benchmark to enhance LLM reasoning on graphs through iterative steps of reasoning, interaction, and execution. At each iteration, the LLM identifies the required information and formulates graph interactions to retrieve relevant data, such as nodes or relationships. These interactions are executed on the graph, and the results are fed back into the reasoning process. This cycle repeats until the LLM concludes the reasoning task and produces the final answer. By iteratively refining its understanding of graph structures, GRAPH-COT effectively handles multi-hop reasoning tasks, ensuring accurate and explainable results in graph-based question answering.
@@ 6.4. Methods Applying the Tree Pipeline
EWEK-QA (ref:ewek-qa, ) triggers two pipelines simultaneously to extract informative knowledge from the web content and the knowledge base. For the retrieval from the web text, EWEK-QA first retrieves the relevant content utilizing the Bing search and then conducts a series of processes to filter and rerank the candidate texts from the web pages. Simultaneously, EWEK-QA avoids retrieval techniques relying on LLMs such as TOG (ref:tog, ) and directly employs a match of representation similarity (cosine score) to obtain the relevant triples. To this end, EWEK-QA further utilizes a pre-trained LLMs to integrate KG triples and web quotes together and construct a unified prompt. By substituting the iterative retrieval in the knowledge base and enhancing it with web search, EWEK-QA significantly improves the quality of extracted knowledge in terms of relevance to the queries without hampering efficiency.
SURGE (ref:SURGE, ) effectively leverages a tree pipeline by structuring its retrieval and generation processes into multiple parallel branches, allowing different retrieval and prompting strategies to execute simultaneously. Instead of sequentially retrieving a single set of facts, SURGE retrieves multiple context-relevant subgraphs in parallel, each representing different aspects of the dialogue history. These subgraphs are processed simultaneously through distinct embedding pathways, ensuring diverse knowledge representations. The model then applies its invariant graph encoding technique to each retrieved subgraph independently, maintaining permutation and relation-inversion invariance across multiple branches. Additionally, during response generation, SURGE integrates these parallel knowledge streams using graph-text contrastive learning, ensuring that the generated response remains consistent across all retrieved knowledge sources. This parallel execution not only enhances efficiency but also enables the model to synthesize information from multiple perspectives, improving response accuracy and informativeness in knowledge-grounded dialogue systems.
@@ 6.5. Comparison between Different Pipelines
Different pipelines offer varying degrees of flexibility and efficiency, influencing how information flows through RAG systems. Sequential pipelines provide a straightforward, stepwise execution, ensuring clarity and simplicity but limiting adaptability when intermediate refinements are needed. In contrast, loop pipelines introduce feedback mechanisms, allowing iterative refinement of retrieved knowledge or prompt modifications, making them well-suited for tasks requiring multiple rounds of reasoning. However, they can introduce computational overhead due to repeated processing. Tree pipelines, on the other hand, enable parallel execution of different retrieval or prompting strategies, improving efficiency in handling diverse queries but potentially increasing system complexity. A future trend of pipelines is moving toward adaptive pipeline selection, where the system dynamically determines the optimal pipeline structure based on the query’s complexity and reasoning needs (jiang2025d, ). This strategy can enable more efficient, context-aware retrieval and generation while minimizing unnecessary computational overhead.
@@ 7.1.1. A Unified View
Knowledge Graph Question Answering (KGQA) targets answering natural language queries by leveraging the structured information encoded in KGs. Unlike traditional question-answering methods that rely solely on text-based retrieval, KGQA focuses on grounding its reasoning in the relationships and entities represented within a knowledge graph. By utilizing the rich semantic structure of KGs, which capture entities, their attributes, and the interrelationships between them, KGQA systems aim to provide precise and contextually accurate answers. This task not only requires understanding the natural language query but also necessitates navigating the graph’s structure to extract relevant facts, perform logical inferences, and deliver a well-supported response. In the ensuing section, we offer in-depth discussions of select noteworthy works, and a more extensive overview of additional works can be found in Figure 5.
@@ 7.1.2. Methods Targeting the KGQA Tasks
G-Retriever (ref:g-retriever, ) is designed to improve KGQA by addressing challenges in reasoning, scalability, and accuracy for real-world textual graphs. It introduces a new benchmark of graph-based question answering, enabling models to handle complex, multi-domain questions beyond basic graph reasoning tasks. A key innovation of this benchmark is formulating the retrieval of subgraph as a Prize-Collecting Steiner Tree optimization (bienstock1993note, ) problem, ensuring relevant information is extracted while maintaining explainability. G-Retriever integrates GNNs and LLMs for fine-tuned reasoning and achieves superior performance across domains such as knowledge graphs, scene graphs, and common-sense reasoning. By offering a conversational interface for graph queries, it advances the explainability, efficiency, and usability of KGQA tasks. This makes it a benchmark in advancing graph-based question-answering systems.
SR (ref:sr, ) aims to enhance KGQA by efficiently retrieving high-quality subgraphs closely aligned with the query. It achieves this by decoupling the retrieval process from the reasoning step, enabling a plug-and-play framework that can be integrated with various KGQA models. SR uses a dual-encoder architecture to expand and refine paths, automatically stopping when relevant subgraphs are constructed. This separation reduces reasoning bias caused by partial subgraphs and improves retrieval precision. By employing weakly supervised and end-to-end fine-tuning strategies, SR demonstrates significant improvements in reasoning accuracy, explainability, and retrieval efficiency, achieving state-of-the-art results in complex multi-hop KGQA tasks.
GraphRAG (ref:graphrag, ) enhances KGQA by integrating traditional RAG with graph-based summarization, addressing global queries that traditional RAG struggles to resolve. It builds an LLM-derived knowledge graph that uses community detection algorithms, such as Leiden, to group related entities and relationships into modular communities. These communities are pre-summarized, allowing for efficient multi-hop reasoning and improved retrieval accuracy. For answering queries, partial responses from community summaries are combined into a global answer through a map-reduce approach. This method improves comprehensiveness, diversity, and scalability, enabling precise, contextually rich responses for large-scale datasets with significantly reduced computational overhead.
Table 5. Summary of different tasks.
@@
KGQA Tasks HSGE (ref:hsge, ), ReTraCk (ref:retrack, ), RNG-KBQA (ref:rng, ), ArcaneQA (ref:arcaneqa, ), EWEK-QA (ref:ewek-qa, ), HamQA (ref:hamqa, ), KG-FiD (ref:kgfid, ), Golden-Retriever (ref:golden, ), SURGE (ref:SURGE, ),KELP (ref:kelp, ), Fact (ref:fact, ), TOG2 (ref:tog2, ), MINERVA (ref:minerva, ), KALMV (ref:kalmv, ),REANO (ref:reano, ),LLaGA (ref:llaga, ), GNN-RAG (ref:gnn-rag, ), G-Retriever (ref:g-retriever, ), GraphRAG (ref:graphrag, ), ROG (ref:rog, ), SR (ref:sr, ), RRA (ref:rra, ), KnowledgeNavigator (ref:kn, ), KG-GPT (ref:kg-gpt, ), KD-COT (ref:kd-cot, ), QA-GNN (ref:qa-gnn, ), MVP-Tuning (ref:mvp-tuning, ), StructGPT (ref:structgpt, ), MHKG (ref:feng, ), MHGRN (ref:noname3, ), Temple-MQA (ref:temple-mqa, ), KagNet (ref:kagnet, ), KG-Agent (ref:kg-agent, ), GenKGQA (ref:gentkgqa, ), GLBK (ref:noname5, ), Keqing (ref:keqing, ), UniKGQA (ref:unikgqa, ), GRAG (ref:grag, ), Knowledge Solver (ref:ksl, ), WebQSQ (ref:webqsp, ), MetaQA (ref:metaqa, ), PullNet (ref:pullnet, ), KnowledGPT (ref:knowledgpt, ), ETD (ref:etd, ), DecAF (ref:decaf, ),HippoRAG (ref:hipporag, ), GGE (ref:gge, ), NSM (ref:nsm, ), SKP (ref:skp, ), KnowGPT (ref:knowgpt, ), Difar (ref:difar, ), Kaping (ref:kaping, ), NuTrea (ref:nutrea, ), OreoLM (ref:oreolm, )
Graph Tasks GPT4Graph (ref:GPT4Graph, ), GraphText (ref:graphtext, ), LLaGA (ref:llaga, ), GRAPH-COT (ref:graphcot, ), Engine (ref:engine, ), GraphBridge (ref:graphbridge, ), MMGCN (ref:mmgcn, ),ConvE (ref:wn, ), Graph-LLM (ref:Graph-LLM, ), GraphGPT (ref:GraphGPT, ), NLGraph (ref:nlgraph, ), GLBench (ref:GLBench, ), GraphEval2000 (ref:GraphEval2000, ), MuseGraph (ref:musegraph, ), Walklm (ref:walklm, )
Domain-specific Tasks GLBK (ref:noname5, ), ATLANTIC (ref:atlantic, ), HyKGE (ref:hykge, ), KG-Rank (ref:kg-rank, ), DALK (ref:dalk, ), KGP (ref:kgp, ), MindMap (ref:mindmap, ), FABULA (ref:fabula, ), FoodGPT (ref:foodgpt, ), KAM-CoT (ref:kamcot, ), MufassirQAS (alan2024rag, ), REALM (zhu2024realm, ), MEDQA (ref:medqa, ), HybridRAG (ref:hybridrag, ), MedGraphRAG (ref:medgraphrag, ), DepsRAG (ref:depsrag, )

### v2

@@ 6.3. Methods Applying the Sequential Pipeline
Keqing (Wang et al., 2023a) adopts a sequential pipeline to address complex multi-hop question answering. It decomposes complex queries into simpler sub-questions using predefined templates. These sub-questions are then aligned with logical chains on a knowledge graph (KG), guiding the retrieval of candidate entities through multi-hop reasoning. A candidate reasoning module evaluates and selects the correct entities from the KG. Finally, the system generates responses by aggregating reasoning paths and answers to provide an interpretable and precise answer to multi-hop questions.
QA-GNN (Yasunaga et al., 2021) integrates LMs and KGs for question answering with a sequential pipeline. The process begins by encoding the QA context including the question and answer options with an LM and retrieving a KG subgraph relevant to the question. A joint ”working graph” is constructed by connecting the QA context node to entities within the KG subgraph. To enhance reasoning, the KG nodes are scored for relevance based on their alignment with the QA context using LM-based scoring. An attention-based GNN performs iterative message passing on this working graph, updating both KG entities and QA context representations. Finally, predictions are made by aggregating the LM output, updated node features, and the working graph representation, achieving accurate and explainable results.
DALK (Li et al., 2024a) makes LLMs and KGs benefit from each other through an elaborate sequential pipeline design. At the beginning of the pipeline, LLMs process the scientific corpus and construct two domain-specific KGs from the related literature via pair-wised relation extraction and generative relation extraction, respectively. This part of the pipeline can be seen as LLMs benefit KGs. Then, DALK utilizes a coarse-to-fine sampling method and a retrieval approach to select knowledge from KGs. The selected knowledge as well as the domain questions are sent to LLMs to generate accurate answers, therefore realizing that KGs benefit LLMs.
@@ 6.4. Methods Applying the Loop Pipeline
RGNN-Ret (Li et al., 2024f) tackles multi-hop reasoning questions by iteratively decomposing the reasoning process into manageable steps. It incorporates a self-critique mechanism that prompts LLMs to generate sub-questions for each reasoning step. After answering a sub-question with retrieved passages, the LLM evaluates whether the accumulated evidence is sufficient to generate a final answer or if further reasoning steps are required. This iterative process enables the system to dynamically adapt to the complexity of the query. To enhance the retrieval quality across reasoning steps, a Recurrent Graph Neural Network (RGNN) is used. The RGNN integrates the graphs of passages from previous reasoning steps, establishing connections between retrieved passages to ensure that relationships between sub-questions and evidence are maintained. By combining semantic distances and contextual information across steps, the RGNN reduces the impact of incorrect sub-questions and improves retrieval accuracy for supporting passages. The iterative loop continues until the self-critique determines that enough evidence has been collected to answer the question.
GRAPH-COT (Jin et al., 2024b) provides a benchmark to enhance LLM reasoning on graphs through iterative steps of reasoning, interaction, and execution. At each iteration, the LLM identifies the required information and formulates graph interactions to retrieve relevant data, such as nodes or relationships. These interactions are executed on the graph, and the results are fed back into the reasoning process. This cycle repeats until the LLM concludes the reasoning task and produces the final answer. By iteratively refining its understanding of graph structures, GRAPH-COT effectively handles multi-hop reasoning tasks, ensuring accurate and explainable results in graph-based question answering.
@@ 6.5. Methods Applying the Tree Pipeline
SURGE (Kang et al., 2022) effectively leverages a tree pipeline by structuring its retrieval and generation processes into multiple parallel branches, allowing different retrieval and prompting strategies to execute simultaneously. Instead of sequentially retrieving a single set of facts, SURGE retrieves multiple context-relevant subgraphs in parallel, each representing different aspects of the dialogue history. These subgraphs are processed simultaneously through distinct embedding pathways, ensuring diverse knowledge representations. The model then applies its invariant graph encoding technique to each retrieved subgraph independently, maintaining permutation and relation-inversion invariance across multiple branches. Additionally, during response generation, SURGE integrates these parallel knowledge streams using graph-text contrastive learning, ensuring that the generated response remains consistent across all retrieved knowledge sources. This parallel execution not only enhances efficiency but also enables the model to synthesize information from multiple perspectives, improving response accuracy and informativeness in knowledge-grounded dialogue systems.
GoT (Besta et al., 2024a) effectively leverages a tree pipeline by expanding an initial prompt into multiple parallel branches of “thoughts,” each branch pursuing a distinct partial solution under a predefined Graph of Operations. Rather than committing to a single chain, GoT repeatedly applies a generate–score–select routine: candidate branches are spawned with a controllable branching factor, locally validated and scored, and then pruned via a keep-best policy to preserve only the most promising subtrees. This staged branching decomposes the task into tractable subproblems that can be solved concurrently within separate branches, while refinement loops allow individual branches to self-correct without restarting the whole pipeline. Crucially, GoT then fuses the surviving branches using explicit aggregation operators, turning the pure tree into a directed acyclic workflow that consolidates partial results into a coherent final output. By combining parallel expansion, principled pruning, iterative refinement, and late-stage aggregation, GoT maintains the efficiency benefits of a tree pipeline across the entire process while achieving higher solution quality and better cost–latency trade-offs than sequential methods.
Beyond-CoT (Yao et al., 2023) leverages a tree-style pipeline by expanding a single query into multiple concurrent branches that persist across two stages—rationale generation and answer inference. First, an Extract–Cluster–Coreference (ECC) routine builds a branched thought graph from the input and optional image caption, effectively fanning out the problem into many node-level sub-traces. Each branch is then encoded in parallel: text via a Transformer encoder, the thought graph via a graph attention network, and vision via a dedicated encoder. Cross-attention aligns tokens with graph nodes, and a gated-fusion layer aggregates these parallel branches into a unified representation that the decoder uses to produce rationales. In stage two, the pipeline re-branches by concatenating the predicted rationales back to the input, reconstructing the thought graph, and repeating the same parallel encoders, alignment, and gated fusion before decoding the final answer. This end-to-end design preserves the efficiency and diversity benefits of a tree pipeline—branching for exploration, parallel encoding for coverage, and late fusion for consolidation—while generalizing the tree into a graph to capture non-linear dependencies.
@@ 6.6. Comparison between Different Pipelines
Different pipelines offer varying degrees of flexibility and efficiency, influencing how information flows through RAG systems. Sequential pipelines provide a straightforward, stepwise execution, ensuring clarity and simplicity but limiting adaptability when intermediate refinements are needed. In contrast, loop pipelines introduce feedback mechanisms, allowing iterative refinement of retrieved knowledge or prompt modifications, making them well-suited for tasks requiring multiple rounds of reasoning. However, they can introduce computational overhead due to repeated processing. Tree pipelines, on the other hand, enable parallel execution of different retrieval or prompting strategies, improving efficiency in handling diverse queries but potentially increasing system complexity. A future trend of pipelines is moving toward adaptive pipeline selection, where the system dynamically determines the optimal pipeline structure based on the query’s complexity and reasoning needs (Jiang et al., 2025). This strategy can enable more efficient, context-aware retrieval and generation while minimizing unnecessary computational overhead.
@@ 7.1.1. A Unified View
Knowledge Graph Question Answering (KGQA) targets answering natural language queries by leveraging the structured information encoded in KGs. Unlike traditional question-answering methods that rely solely on text-based retrieval, KGQA focuses on grounding its reasoning in the relationships and entities represented within a knowledge graph. By utilizing the rich semantic structure of KGs, which capture entities, their attributes, and the interrelationships between them, KGQA systems aim to provide precise and contextually accurate answers. This task not only requires understanding the natural language query but also necessitates navigating the graph’s structure to extract relevant facts, perform logical inferences, and deliver a well-supported response. In the ensuing section, we offer in-depth discussions of select noteworthy works, and a more extensive overview of additional works can be found in Figure 5. To better summarize these methods, we leave the brief introduction of them in the Appendix.
@@ 7.1.2. Methods Targeting the KGQA Tasks
G-Retriever (He et al., 2024) is designed to improve KGQA by addressing challenges in reasoning, scalability, and accuracy for real-world textual graphs. It introduces a new benchmark of graph-based question answering, enabling models to handle complex, multi-domain questions beyond basic graph reasoning tasks. A key innovation of this benchmark is formulating the retrieval of subgraph as a Prize-Collecting Steiner Tree optimization (Bienstock et al., 1993) problem, ensuring relevant information is extracted while maintaining explainability. G-Retriever integrates GNNs and LLMs for fine-tuned reasoning and achieves superior performance across domains such as knowledge graphs, scene graphs, and common-sense reasoning. By offering a conversational interface for graph queries, it advances the explainability, efficiency, and usability of KGQA tasks. This makes it a benchmark in advancing graph-based question-answering systems.
GraphRAG (Edge et al., 2024) enhances KGQA by integrating traditional RAG with graph-based summarization, addressing global queries that traditional RAG struggles to resolve. It builds an LLM-derived knowledge graph that uses community detection algorithms, such as Leiden, to group related entities and relationships into modular communities. These communities are pre-summarized, allowing for efficient multi-hop reasoning and improved retrieval accuracy. For answering queries, partial responses from community summaries are combined into a global answer through a map-reduce approach. This method improves comprehensiveness, diversity, and scalability, enabling precise, contextually rich responses for large-scale datasets with significantly reduced computational overhead.
Table 5. Summary of different tasks. The brief introduction of the listed methods can be found in the appendix.
@@
KGQA Tasks HSGE (Sun et al., 2023), ReTraCk (Chen et al., 2021), RNG-KBQA (Ye et al., 2021), ArcaneQA (Gu and Su, 2022), EWEK-QA (Dehghan et al., 2024), HamQA (Dong et al., 2023b), KG-FiD (Yu et al., 2022), Golden-Retriever (An et al., 2024), SURGE (Kang et al., 2022),KELP (Liu et al., 2024b), Fact (Opsahl, 2024), TOG2 (Ma et al., 2024), MINERVA (Das et al., 2018), KALMV (Baek et al., 2023c),REANO (Fang et al., 2024),LLaGA (Chen et al., 2024b), GNN-RAG (Mavromatis and Karypis, 2024), G-Retriever (He et al., 2024), GraphRAG (Edge et al., 2024), ROG (Luo et al., 2024b), SR (Zhang et al., 2022), RRA (Wu et al., 2023b), KnowledgeNavigator (Guo et al., 2024), KG-GPT (Kim et al., 2023a), KD-COT (Wang et al., 2023b), QA-GNN (Yasunaga et al., 2021), MVP-Tuning (Huang et al., 2023b), StructGPT (Jiang et al., 2023a), MHKG (Feng et al., 2020), MHGRN (Chakraborty, 2024), Temple-MQA (Cheng et al., 2024), KagNet (Lin et al., 2019), KG-Agent (Jiang et al., 2024a), GenKGQA (Gao et al., 2024b), GLBK (Delile et al., 2024), Keqing (Wang et al., 2023a), UniKGQA (Jiang et al., 2023b), GRAG (Hu et al., 2024), Knowledge Solver (Feng et al., 2023), WebQSQ (Yih et al., 2016), MetaQA (Zhang et al., 2018), PullNet (Sun et al., 2019), KnowledGPT (Wang et al., 2023c), ETD (Liu et al., 2024a), DecAF (Yu et al., 2023),HippoRAG (Gutiérrez et al., 2024), GGE (Gao et al., 2022), NSM (He et al., 2021), SKP (Dong et al., 2023a), KnowGPT (Zhang et al., 2024c), Difar (Baek et al., 2023a), Kaping (Baek et al., 2023b), NuTrea (Choi et al., 2023), OreoLM (Hu et al., 2022)
Graph Tasks GPT4Graph (Guo et al., 2023), GraphText (Zhao et al., 2023a), LLaGA (Chen et al., 2024b), GRAPH-COT (Jin et al., 2024b), Engine (Zhu et al., 2024c), GraphBridge (Wang et al., 2024c), MMGCN (Wei et al., 2019),ConvE (Dettmers et al., 2018), Graph-LLM (Chen et al., 2024c), GraphGPT (Tang et al., 2024), NLGraph (Wang et al., 2024a), GLBench (Li et al., 2024d), GraphEval2000 (Wu et al., 2024c), MuseGraph (Tan et al., 2024a), Walklm (Tan et al., 2024b)
Domain-specific Tasks GLBK (Delile et al., 2024), ATLANTIC (Munikoti et al., 2023), HyKGE (Jiang et al., 2024c), KG-Rank (Yang et al., 2024a), DALK (Li et al., 2024a), KGP (Wang et al., 2024d), MindMap (Wen et al., 2024), FABULA (Ranade and Joshi, 2023), FoodGPT (Qi et al., 2023), KAM-CoT (Mondal et al., 2024), MufassirQAS (Alan et al., 2024), REALM (Zhu et al., 2024b), MEDQA (Jin et al., 2020), HybridRAG (Sarmah et al., 2024), MedGraphRAG (Wu et al., 2024b), DepsRAG (Alhanahnah et al., 2024)